A 100.0 mL sample of 0.300 M NaOH is mixed with a 100.0 mL sample of 0.300 M HNO3 in a coffee cup calorimeter. If both solutions were initially at 35.0°C and the temperature of the resulting solution was recorded as 37.0°C, determine the DH°rxn (in units of kJ/mol NaOH) for the neutralization reaction between aqueous NaOH and HCl. Assume 1) that no heat is lost to the calorimeter or the surroundings, and 2)that the density(1.00 g/mL) and the specific heat capacity (4.18 J/g oC) of the resulting solution are the same as water.

A.-16.7 kJ/mol NaOH

B. -27.9 kJ/mol NaOH

C. -169 kJ/mol NaOH

D. -34.4 kJ/mol NaOH

E. -55.7 kJ/mol NaOH
Answer #1

Since temperature of the solution rises, the reaction is exothermic in nature and therefore ∆H° is negative.

ULUI. * Total volume o solution = (100.0 + 100:o) mL = 200.0 ml Mass of solution = v*d = 200.0 ml x 1.00 g Now, heat change,